If Tom is reading a book that has 193 pages and he has read 3 pages more than one forth of the number of pages he hasn't read.
How many pages does he have left to read?

Answer by nerdybill(7384)

If Tom is reading a book that has 193 pages and he has read 3 pages more than one forth of the number of pages he hasn't read.
.
Let x = number of pages he hasn't read
then
(1/4)x + 3 = number of pages read
.
x + (1/4)x + 3 =193
4x + x + 12 = 772
5x + 12 = 772
5x = 760
x = 152
.
How many pages does he have left to read?
152